# Designing Human-AI Systems: Anthropomorphism and Framing Bias on Human-AI Collaboration

### v1

@@ Abstract
AI is redefining how humans interact with technology, leading to a synergetic collaboration between the two. Nevertheless, the effects of human cognition on this collaboration remain unclear. This study investigates the implications of two cognitive biases, anthropomorphism and framing effect, on human-AI collaboration within a hiring setting. Subjects were asked to select job candidates with the help of an AI-powered recommendation tool. The tool was manipulated to have either human-like or robot-like characteristics and presented its recommendations in either positive or negative frames. The results revealed that the framing of AI’s recommendations had no significant influence on subjects’ decisions. In contrast, anthropomorphism significantly affected subjects’ agreement with AI recommendations. Contrary to expectations, subjects were less likely to agree with the AI if it had human-like characteristics. These findings demonstrate that cognitive biases can impact human-AI collaboration and highlight the need for tailored approaches to AI product design, rather than a single, universal solution.
@@ 1 Introduction
This study investigates the impact of two cognitive biases, anthropomorphism and framing effect, on human-AI collaboration in the context of hiring decision-making. An experiment was designed to simulate the screening phase of the recruitment process, where companies are already using AI-based tools (see Section 2.2 and Section 2.3). This study aims to expand our understanding of the cognitive consequences of human-AI collaboration and contribute to the design of AI-based products that take into account these cognitive factors, enhancing AI efficiency while also considering its ethical implications.
@@ 2 Related work
To ensure that AI can comprehend human interpretation, it is essential to integrate human cognition into the development of AI systems. Human biases are a crucial component of human cognition, yet research on their implications in human-AI collaboration is scarce \@BBOPcitep\@BAP\@BBN(Rastogi et al.,, 2022)\@BBCP. I aim to shed light on cognitive biases and their consequences in human-AI collaboration, thereby contributing to the ongoing discussion. The present study explores two different ways in which we can design AI tools in a way that enhances human-AI collaboration by taking into account human cognition and its biases.
@@ 2.1 Cognitive bias and dual process theories
While the study of heuristics dates back several centuries, the work of Daniel Kahneman and Amos Tversky, inspired by Herbert Simon’s ideas, marked the beginning of a new era in the study of biases in human judgment and decision-making \@BBOPcitep\@BAP\@BBN(Hjeij and Vilks,, 2023)\@BBCP. However, the concept of heuristics is still controversial \@BBOPcitep\@BAP\@BBN(Neth and Gigerenzer,, 2015)\@BBCP. In this work, I adopt Kahneman and Tversky’s definition of heuristics as mental methods that help humans make quick judgments under uncertainty. These heuristics often lead to systematic errors, which are referred to as cognitive biases \@BBOPcitep\@BAP\@BBN(Tversky and Kahneman,, 1974)\@BBCP.
@@ 2.4 Cognitive biases and Human-AI Collaboration
This work examines how an AI decision-aid tool might influence the recruiter’s decision-making process when AI’s recommendations show two types of biases: anthropomorphism and the framing effect. This study should provide us with more information about the effect of cognitive biases on human-AI interaction. Ultimately, I am interested in testing the following two hypotheses:
    H1: Framing of the AI recommendation can alter the degree to which the human conforms to it (framing effect).
@@ 2.4.1 Anthropomorphism and trust
By anthropomorphism, we refer to the tendency to ascribe human traits to non-human entities, in the context of social interactions, to rationalize their behavior \@BBOPcitep\@BAP\@BBN(Duffy,, 2003)\@BBCP. Anthropomorphism can be considered a cognitive bias since it consists of an erroneous human perception of inanimate beings caused by an unconscious process \@BBOPcitep\@BAP\@BBN(Dacey,, 2017)\@BBCP. That is to say, it is associated with the heuristic and automatic type 1 processes. This human phenomenon is caused by a set of human-like features called anthropomorphic design \@BBOPcitep\@BAP\@BBN(Lemaignan et al.,, 2014)\@BBCP.
Although anthropomorphism can be categorized as a cognitive bias (i.e., a systematic error), it can have a positive impact when applied to AI tools. Indeed, a compelling reason for adopting anthropomorphic approaches is their potential to facilitate human-AI interaction \@BBOPcitep\@BAP\@BBN(Duffy,, 2003)\@BBCP. One way in which anthropomorphic design can improve interaction is by increasing human trust in AI. According to previous research, the extent to which humans are willing to depend on automated systems in uncertain scenarios is largely influenced by trust \@BBOPcitep\@BAP\@BBN(Hoff and Bashir,, 2015)\@BBCP. Moreover, trust plays a critical role in mitigating the uncertainty associated with the use of AI among humans, thereby determining the acceptance and adoption of such technologies \@BBOPcitep\@BAP\@BBN(Lukyanenko et al.,, 2022)\@BBCP.
@@
Regardless of the aforementioned cases, few studies have investigated how anthropomorphism can affect human trust in automatic systems that assist in decision-making \@BBOPcitep\@BAP\@BBN(Pak et al.,, 2012)\@BBCP. Additionally, the literature on how the anthropomorphic design of AI influences user trust remains scarce \@BBOPcitep\@BAP\@BBN(Troshani et al.,, 2021)\@BBCP. I could not find any research that studied how the appearance of AI applications that assist hiring managers in the recruitment process can affect their trust, even though they are commonly used by companies.
Non-significant correlations between an anthropomorphic design and human trust may suggest that subjects are already anthropomorphizing the AI during its interactions, as some previous studies suggest \@BBOPcitep\@BAP\@BBN(Chinmulgund et al.,, 2023)\@BBCP. Likewise, a negative correlation could imply that the attribution of human-like characteristics to an AI agent is actually decreasing human affinity, which might be related to the uncanny valley effect (see \@BBOPcitep\@BAP\@BBNMori,, 1970\@BBCP). A positive correlation, on the other hand, might indicate that providing human traits to an AI makes subjects view the AI as more human.
@@ 2.4.2 Framing effect
The concept of “framing effect” refers to the tendency of decision-makers to respond differently when presented with distinct (yet equivalent) descriptions of the same problem \@BBOPcitep\@BAP\@BBN(Frisch,, 1993)\@BBCP. Although framing effects tend to be treated homogeneously, we can distinguish three different types of framing effects: risky choice framing, goal framing, and attribute framing \@BBOPcitep\@BAP\@BBN(Levin et al.,, 1998)\@BBCP. Risky choice framing has often been regarded as the prototypical framing effect due to the repercussion of the “Asian disease problem”, which studied risky choice framing and was introduced by \@BBOPcitet\@BAP\@BBNTversky and Kahneman, (1981)\@BBCP. Nevertheless, both goal framing and attribute framing have also been extensively studied in the literature \@BBOPcitep\@BAP\@BBN(Levin et al.,, 1998)\@BBCP.
@@
The present study focuses on attribute framing. Attribute framing addresses how the description of a single attribute of an object can impact decision-maker preferences \@BBOPcitep\@BAP\@BBN(Levin et al.,, 1998)\@BBCP. The attribute is typically described in two different ways, using a positive or a negative frame. The evidence indicates that individuals evaluate the object more favorably when the positive frame is presented to them \@BBOPcitep\@BAP\@BBN(Leong et al.,, 2017)\@BBCP. For example, a study found out that job candidates who are exposed during realistic job previews to positively framed messages (e.g., “Eighty-five percent of the employees who do this job are satisfied with the job”) have higher job expectations and willingness to accept a job offer than those that were given the same information with a negative frame (e.g., “Fifteen percent of the employees who do this job are dissatisfied with the job”), \@BBOPcitep\@BAP\@BBN(Buda and Charnov,, 2003)\@BBCP.
@@
In a recent study, \@BBOPcitet\@BAP\@BBNJin et al., (2017)\@BBCP examined how online consumers modify their purchasing decisions depending on the framing of product attribute descriptions. Their results showed that consumers under the positive framing condition had a higher purchasing intention compared to the negative framing condition. Moreover, negative messages lead to increased cognitive conflict and difficulty in decision-making. These findings could have practical implications for the development of AI-based products. For instance, AI chatbots on websites might be instructed by companies to present the product’s attributes using positive frames.
@@
Nevertheless, the use of AI tools also raises ethical concerns, particularly in certain domains, such as the recruitment sector. Inconsistent use of positive and negative frames in AI screening tools may lead to candidate discrimination. AI-enabled recruiting tools are already adopted by organizations, and due to their current design, recruiters may be subject to framing effects. AI-powered screening software [^1] See for example Ideal, a tech company acquired by Ceridian which developed an AI tool that allows recruiters to “instantly screen and shortlist thousands of candidates”. typically grades and ranks candidates based on their predicted fit for the role. One of the aspects emphasized by these tools is, for instance, the relative position of a specific candidate within the pool of applicants (e.g., “There is a 90% probability of hiring someone worse than this candidate”). Note that the same information could be framed differently. It would be equivalent to saying that “there is a 10% probability of hiring someone as good or better than this candidate”.
@@
Previous research on the framing effect suggests that the frame used by the AI tool could potentially affect the recruiter’s perception of the candidate. If this is the case, it would be desirable to design AI tools where the framing effect is taken into account in order to avoid candidates’ discrimination. To the extent that humans are susceptible to cognitive biases, the decision-aid technologies we develop should be designed in a way that mitigates these biases \@BBOPcitep\@BAP\@BBN(Arkes,, 1991)\@BBCP. For example, ensuring that AI tools display two different frames or the same frame for all the different candidates.
@@ 3 Experiment
For this study, I designed a binary prediction task within a recruitment setting to address two main research questions. First, I build on the work of \@BBOPcitep\@BAP\@BBN(Rastogi et al.,, 2022)\@BBCP and investigate the impact of cognitive bias in the context of AI-assisted decision-making. Particularly, I focus on attribute framing by measuring the effect of positively and negatively framed AI recommendations. Second, I aim to understand how AI appearance (either human-like or robot-like) affects human trust, following previous work on the subject \@BBOPcitep\@BAP\@BBN(Pak et al.,, 2012)\@BBCP.
@@ 3.1 Participants
To run the experiment, 750 participants from the Prolific platform were recruited. Since the survey was in English, the pool was limited to English native speakers from the United States and the United Kingdom. Participants were compensated £7.5 per hour, which the platform regarded as fair.
@@ 3.2 Task and AI model
The task aimed to emulate the screening phase of a recruitment process within a private company. Participants were asked to act as recruiters and decide whether a student should be invited for an interview or not. To help them in their decisions, I provided some details about the candidates (e.g., grades, work experience, area of specialization), and an AI recommendation.
The dataset used provided information about 215 candidates, including the aforementioned details and their employment situation (“placed” or “not placed”). The dataset and the methodology employed to generate the AI recommendations were directly sourced from the discussions and resources available on the Campus Recruitment contest on Kaggle.com \@BBOPcitep\@BAP\@BBN(Dhimant,, 2020)\@BBCP. After handling missing values, I preprocessed the data by encoding the categorical variables into numerical values and splitting the data into training and testing sets (80:20 split). Then, I fit a logistic regression model on the dataset, obtaining an accuracy rate of 83%. Additionally, I added a new column with the predicted probability for each candidate within the test set. These predicted probabilities were used for the AI recommendation and were unique per candidate. However, the form in which they were presented to subjects varied, as will be discussed in another section.
To create a more realistic scenario, I provided a brief description of the company and the position in the instructions part, before the actual survey. Furthermore, a name was attributed to each participant. To avoid possible discriminations that could invalidate the results, I tried to select names as neutral as possible. An equal distribution between names of diverse gender identities was also pursued.
3.3 Survey
The survey was made on Qualtrics, and it had fifteen questions. Every question presented a distinct candidate with different characteristics. In each of these questions, subjects were asked to decide whether to invite the candidate to the interview or not. The questions had identical structure and predictive tasks. They included a section on the academic details, another one on the employability details, and a final one with the AI recommendation (see Figure 2). This last section was the only one that varied among subjects and will be explained in a separate part.
@@ 3.3 Survey
To emulate the recruiter’s prior experience and enhance the realism of the survey, I included a training section consisting of five questions. Candidates in this section were identical for all subjects, but they were displayed in a random order. The candidates were selected so that they covered all the ranges from [0.5, 0.6] to (0.9, 1] of the model’s estimated probability. Using these estimations as a proxy for difficulty, this selection ensured that the training section contained diverse levels of difficulty \@BBOPcitep\@BAP\@BBN(Rastogi et al.,, 2022)\@BBCP.
@@
Figure 3: Example of one of the two attention checks. Although the structure of the slide remains the same, the question changes. Instead of the question: “Do you invite the candidate for an interview”, subjects are now asked “Did Corina specialize in Arts during Higher Secondary Education?”.

### v2

@@ Abstract
AI is redefining how humans interact with technology, leading to a synergetic collaboration between the two. Nevertheless, the effects of human cognition on this collaboration remain unclear. This study investigates the implications of two cognitive biases, anthropomorphism and framing effect, on human-AI collaboration within a hiring setting. Subjects were asked to select job candidates with the help of an AI-powered recommendation tool. The tool was manipulated in a 3 x 3 between-subjects design to present three different AI identities (human-like, robot-like, generic) and three types of framing (positive, negative, and neutral). The results revealed that the framing of AI’s recommendations had no significant influence on subjects’ decisions. In contrast, anthropomorphism significantly affected subjects’ agreement with AI recommendations. Subjects were less likely to agree with the AI if it had human-like characteristics. These findings demonstrate that cognitive biases can impact human-AI collaboration and highlight the need for tailored approaches to AI product design, rather than a single, universal solution.
Keywords: artificial intelligence; human-AI collaboration; human-centered AI; cognitive biases; decision support systems.
@@ 1 Introduction
This study investigates the impact of two cognitive biases, anthropomorphism and framing effect, on human-AI collaboration in the context of hiring decision-making. An experiment was designed to simulate the screening phase of the recruitment process, where companies are already using AI-based tools (see Section 2.2 and Section 2.3). This study aims to expand our understanding of the cognitive consequences of human-AI collaboration and contribute to the design of AI-based products that consider these cognitive factors.
@@ 2 Related work
To ensure that AI can comprehend human interpretation, it is essential to integrate considerations of human cognition into the development of AI systems. Human biases are a crucial component of human cognition, yet research on their implications in human-AI collaboration is scarce \@BBOPcitep\@BAP\@BBN(Rastogi et al.,, 2022)\@BBCP. The present study explores how we should design AI tools by investigating cognitive biases and their consequences in human-AI collaboration.
@@ 2.1 Cognitive bias and dual process theories
While the study of heuristics dates back several centuries, the work of Daniel Kahneman and Amos Tversky, inspired by Herbert Simon’s ideas, marked the beginning of a new era in the study of biases in human judgment and decision-making \@BBOPcitep\@BAP\@BBN(Hjeij and Vilks,, 2023)\@BBCP. However, the concept of heuristics is still controversial \@BBOPcitep\@BAP\@BBN(Neth and Gigerenzer,, 2015)\@BBCP. In this work, we adopt Kahneman and Tversky’s definition of heuristics as mental methods that help humans make quick judgments under uncertainty. These heuristics often lead to errors in decision-making, which are referred to as cognitive biases \@BBOPcitep\@BAP\@BBN(Tversky and Kahneman,, 1974)\@BBCP.
@@ 2.4 Cognitive biases and Human-AI Collaboration
This work examines how an AI decision-aid tool might influence the recruiter’s decision-making process when AI’s recommendations show two types of biases: anthropomorphism and the framing effect. This study should provide us with more information about the effect of cognitive biases on human-AI interaction. Ultimately, it tests the following two hypotheses:
    H1: Framing of the AI recommendation will influence the degree to which the human conforms to it (framing effect).
@@ 2.4.1 Anthropomorphism and trust
By anthropomorphism, we refer to the tendency to ascribe human traits to non-human entities, in the context of social interactions, to rationalize their behavior \@BBOPcitep\@BAP\@BBN(Duffy,, 2003)\@BBCP. Anthropomorphism can be considered a cognitive bias since it consists of an erroneous human perception caused by an unconscious process \@BBOPcitep\@BAP\@BBN(Dacey,, 2017)\@BBCP. That is to say, it is associated with the heuristic and automatic type 1 processes.
Researchers have already explored the potential of intentionally incorporating human-like features (anthropomorphic design) to amplify existing tendencies toward anthropomorphism. This approach aims to enhance human interaction with technology, particularly in fields like human-robot interaction \@BBOPcitep\@BAP\@BBN(Fink,, 2012)\@BBCP. Anthropomorphic designs could also facilitate human-AI interaction in various ways \@BBOPcitep\@BAP\@BBN(Duffy,, 2003)\@BBCP, one of which is by increasing human trust in AI. According to previous research, the extent to which humans are willing to depend on automated systems in uncertain scenarios is largely influenced by trust \@BBOPcitep\@BAP\@BBN(Hoff and Bashir,, 2015)\@BBCP. Moreover, trust plays a critical role in mitigating the uncertainty associated with the use of AI among humans, thereby determining the acceptance and adoption of such technologies \@BBOPcitep\@BAP\@BBN(Lukyanenko et al.,, 2022)\@BBCP.
@@
Regardless of the aforementioned cases, few studies have investigated how anthropomorphism can affect human trust in automatic systems that assist in decision-making \@BBOPcitep\@BAP\@BBN(Pak et al.,, 2012)\@BBCP. Additionally, the literature on how the anthropomorphic design of AI influences user trust remains scarce \@BBOPcitep\@BAP\@BBN(Troshani et al.,, 2021)\@BBCP. Although companies are already using AI tools to assist their hiring processes, no research has been found on how their anthropomorphic design could affect recruiters’ trust.
A non-significant association between anthropomorphic design and human trust may suggest that subjects are already anthropomorphizing the AI during its interactions, as some previous studies suggest \@BBOPcitep\@BAP\@BBN(Chinmulgund et al.,, 2023)\@BBCP. Likewise, a negative association could imply that the attribution of human-like characteristics to an AI agent is actually decreasing human affinity, which might be related to the uncanny valley effect (see \@BBOPcitep\@BAP\@BBNMori,, 1970\@BBCP). On the other hand, a positive association might indicate that providing human traits to an AI makes subjects view the AI as more human.
@@ 2.4.2 Framing effect
The concept of “framing effect” refers to the tendency of decision-makers to respond differently when presented with distinct (yet equivalent) descriptions of the same problem \@BBOPcitep\@BAP\@BBN(Frisch,, 1993)\@BBCP. Although framing effects tend to be treated homogeneously, we can distinguish three different types of framing effects: risky choice framing, goal framing, and attribute framing \@BBOPcitep\@BAP\@BBN(Levin et al.,, 1998)\@BBCP. Risky choice framing has often been regarded as the prototypical framing effect due to the repercussion of the “Asian disease problem”, which was introduced by \@BBOPcitet\@BAP\@BBNTversky and Kahneman, (1981)\@BBCP. In this experiment, subjects shifted from risk aversion to risk-taking attitudes depending on the framing of the outcome. When the outcome of a decision was described by the number of lives saved (positive frame), the majority choice was risk averse. On the contrary, if the outcome was described by the number of lives lost (negative frame), the majority choice was risk-taking. Goal framing measures how subjects’ decisions deviate depending on how the consequences of their actions are framed. In the positively framed condition, only the potential benefits of achieving a specific result are shown to participants whereas in the negatively framed conditions, the focus is shifted to the negative consequences that would be avoided by that result \@BBOPcitep\@BAP\@BBN(Levin et al.,, 1998)\@BBCP.
@@
The present study focuses on attribute framing, which has also been extensively studied in the literature. Attribute framing addresses how the description of a single attribute of an object can impact decision-maker preferences \@BBOPcitep\@BAP\@BBN(Levin et al.,, 1998)\@BBCP. The attribute is typically described in two different ways, using a positive or a negative frame. The evidence indicates that individuals evaluate the object more favorably when the positive frame is presented to them \@BBOPcitep\@BAP\@BBN(Leong et al.,, 2017)\@BBCP. For example, a study found out that job candidates who are exposed during realistic job previews to positively framed messages (e.g., “Eighty-five percent of the employees who do this job are satisfied with the job”) have higher job expectations and willingness to accept a job offer than those that were given the same information with a negative frame (e.g., “Fifteen percent of the employees who do this job are dissatisfied with the job”), \@BBOPcitep\@BAP\@BBN(Buda and Charnov,, 2003)\@BBCP.
@@
In a recent study, \@BBOPcitet\@BAP\@BBNJin et al., (2017)\@BBCP examined how online consumers modify their purchasing decisions depending on the framing of product attribute descriptions. Their results showed that consumers under the positive framing condition had a higher purchasing intention compared to the negative framing condition. Moreover, negative messages led to increased cognitive conflict and difficulty in decision-making. These findings could have practical implications for the development of AI-based products. For instance, AI chatbots on websites might be instructed by companies to present the product’s attributes using positive frames.
@@
Nevertheless, the use of AI tools also raises ethical concerns, particularly in certain domains, such as the recruitment sector. Inconsistent use of positive and negative frames in AI screening tools may lead to candidate discrimination. AI-enabled recruiting tools are already adopted by organizations, and due to their current design, recruiters may be subject to framing effects. AI-powered screening software [^1] See for example Ideal, a tech company acquired by Ceridian which developed an AI tool that allows recruiters to “instantly screen and shortlist thousands of candidates.” typically grades and ranks candidates based on their predicted fit for the role. These tools provide recruiters with the relative position of a specific candidate within the pool of applicants (e.g., “There is a 90% probability of hiring someone worse than this candidate”). Note that the same information could be framed differently. It is equivalent to saying that “there is a 10% probability of hiring someone as good or better than this candidate.”
@@
Previous research on the framing effect suggests that the frame used by an AI tool could potentially affect the recruiter’s perception of the candidate. If this is the case, it would be desirable to design AI tools where the framing effect is taken into account in order to avoid candidates’ discrimination. To the extent that humans are susceptible to cognitive biases, the decision-aid technologies we develop should be designed in a way that mitigates these biases \@BBOPcitep\@BAP\@BBN(Arkes,, 1991)\@BBCP. For example, ensuring that AI tools display two different frames or the same frame for all the different candidates.
@@ 3 Experiment
For this study, we designed a binary prediction task within a recruitment setting to address two main research questions. First, we build on the work of \@BBOPcitep\@BAP\@BBN(Rastogi et al.,, 2022)\@BBCP and investigate the impact of cognitive bias in the context of AI-assisted decision-making. Particularly, we focus on attribute framing by measuring the effect of positively and negatively framed AI recommendations. Second, we aim to understand how AI appearance (either human-like or robot-like) affects human trust, following previous work on the subject \@BBOPcitep\@BAP\@BBN(Pak et al.,, 2012)\@BBCP.
@@ 3.1 Participants
To run the experiment, 750 participants from Prolific were recruited. Prolific is an online platform that allows researchers to recruit participants from a diverse pool. Since the survey was in English, the pool was limited to English native speakers from the United States and the United Kingdom. Participants were compensated £7.5 per hour, which the platform regarded as fair based on several factors, such as the complexity of the task or the amount of effort required.
@@ 3.2 Task and AI model
The task aimed to emulate the screening phase of a recruitment process within a private company. Participants were asked to act as recruiters and decide whether a student should be invited for an interview or not. To help them in their decisions, we provided some details about the candidates (e.g., grades, work experience, area of specialization), and an AI recommendation.
The dataset used provided information about 215 candidates, including the aforementioned details and their employment situation (“placed” or “not placed”). The dataset and the methodology employed to generate the AI recommendations were directly sourced from the discussions and resources available on the Campus Recruitment contest on Kaggle.com \@BBOPcitep\@BAP\@BBN(Dhimant,, 2020)\@BBCP. After handling missing values, we preprocessed the data by encoding the categorical variables into numerical values and splitting the data into training and testing sets (80:20 split). Then, we fit a logistic regression model on the dataset, obtaining an accuracy rate of 83%. Additionally, we added a new column with the predicted probability for each candidate within the test set. These predicted probabilities were used for the AI recommendation and were unique per candidate. However, the form in which they were presented to subjects varied, as will be discussed in another section.
To create a more realistic scenario, we briefly described the company and the position in the instructions part before the actual survey. Furthermore, a name was attributed to each participant. To avoid possible discriminations that could invalidate the results, we tried to select names as neutral as possible. An equal distribution between names of diverse gender identities was also pursued.
3.3 Study Procedure
The survey was made on Qualtrics, and it had fifteen questions. Every question presented a distinct candidate with different characteristics. In each of these questions, subjects were asked to decide whether to invite the candidate to the interview or not. The questions had identical structure and predictive tasks. They included a section on the academic details, another one on the employability details, and a final one with the AI recommendation (see Figure 2). AI’s recommendation section was the only one that changed among subjects.
@@ 3.3 Study Procedure
To emulate the recruiter’s prior experience and enhance the realism of the survey, we included a training section consisting of five questions. Candidates in this section were identical for all subjects, but they were displayed in a random order. The candidates were selected so that they covered all the ranges from [0.5, 0.6] to (0.9, 1] of the model’s estimated probability. Using these estimations as a proxy for difficulty, this selection ensured that the training section contained diverse levels of difficulty \@BBOPcitep\@BAP\@BBN(Rastogi et al.,, 2022)\@BBCP.
@@
Figure 3: Example of one of the two attention checks. Although the structure of the slide remains the same, the question changes. Instead of the question: “Do you invite the candidate for an interview”, subjects are now asked “Did Corina specialize in Arts during Higher Secondary Education?”